OpenStudy (anonymous):

Help with mixture problems? Amy mixes 12 liters of 16% acid solution with a 40% acid solution which results in 32% acid solution. Let x represent the number of liters of 40% solution used. Which equation can be used to find the number of liters of 40% acid solution in the mixture? A. 1.92 + 0.32x = 0.4(x + 12) B. 3.84 + 0.4x = 0.16(x + 12) C. 4.8 + 0.32x = 0.4(x + 12) D. 1.92 + 0.4x = 0.32(x + 12) Please also show me HOW you got your answer. Thanks so much!

OpenStudy (anonymous):

@Directrix @SolomonZelman

OpenStudy (anonymous):

@ganeshie8 @cwrw238

OpenStudy (cwrw238):

the volume of pure acid in 12 liters of 16% strength = 12*0.16 = 1.92 liters and for the 40% solution its 0.40 x where x is the volume required the resulting mixture contains 32% of pure acid and its volume is x + 12 liters so volume of pure acis is 0.32(x + 12) so we have 1.92 + 0.40x = 0.32(x + 12)

OpenStudy (cwrw238):

option D

OpenStudy (anonymous):

Ok, thank you so much!

OpenStudy (cwrw238):

yw